### v1

# Does the 220 PeV Event at KM3NeT Point to New Physics?

Abstract
The KM3NeT collaboration recently reported the observation of KM3-230213A, a neutrino event with an energy exceeding 100 PeV, more than an order of magnitude higher than the most energetic neutrino in IceCube’s catalog. Given its longer data-taking period and larger effective area relative to KM3NeT, IceCube should have observed events around that energy. This tension has recently been quantified to lie between $2\sigma$ and $3.5\sigma$, depending on the neutrino source. A $\mathcal{O}(100)$ PeV neutrino detected at KM3NeT has traversed approximately $147$ km of rock and sea en route to the detector, whereas neutrinos arriving from the same location in the sky would have only traveled through about $14$ km of ice before reaching IceCube. We use this difference in propagation distance to address the tension between KM3NeT and IceCube. Specifically, we consider a scenario in which the source emits sterile neutrinos that partially convert to active neutrinos through oscillations. We scrutinize two such realizations, one where a new physics matter potential induces a resonance in sterile-to-active transitions and another one where off-diagonal neutrino non-standard interactions are employed. In both cases, sterile-to-active neutrino oscillations become relevant at length scales of $\sim 100$ km, resulting in increased active neutrino flux near the KM3NeT detector, alleviating the tension between KM3NeT and IceCube. Overall, we propose the exciting possibility that neutrino telescopes may have started detecting new physics.
Introduction. The field of high-energy neutrino astronomy kicked off in 2013 with the discovery of a number of neutrinos with energies between $100$ TeV and $1$ PeV by IceCube IceCube:2013low ; IceCube:2014stg . This count has steadily increased over the years IceCube:2020wum , including the observation of a Glashow resonance event IceCube:2021rpz . IceCube has also identified several high-energy neutrino sources, including the blazar TXS 0506+056 IceCube:2018cha ; IceCube:2018dnn , NGC 1068 active galaxy IceCube:2022der , and the Milky Way IceCube:2023ame . In parallel with IceCube, located at the South Pole, the ANTARES experiment was taking data in the Mediterranean Sea, using water instead of ice as the detection medium ANTARES:1999fhm ; ANTARES:2011hfw . ANTARES was succeeded by KM3NeT Margiotta:2014gza , which consists of two detectors at different locations: ARCA and ORCA, where the former is specialized for high-energy astrophysical neutrino searches. It was recently reported KM3NeT:2025npi (see also companion papers KM3NeT:2025vut ; KM3NeT:2025bxl ; KM3NeT:2025aps ; KM3NeT:2025ccp ; KM3NeT:2025mfl ) that ARCA detected a neutrino event on 13 February 2023, named KM3-230213A, with median neutrino energy of $E_{\nu}=220$ PeV. This sets the record for the highest-energy neutrino ever observed, exceeding by $\mathcal{O}(10)$ the highest-energy neutrino in IceCube’s dataset.
Comparing the effective areas and data-taking time across KM3NeT and IceCube, it is unexpected that KM3NeT would have been the first experiment observing $\mathcal{O}(100)$ PeV neutrinos. In fact, the observation of a single event at KM3NeT would point to the expectation of at least several events of similar energy reported by IceCube. The statistical significance of this tension was quantified in Li:2025tqf , where the authors obtained results between $2\sigma$ and $3.5\sigma$, depending on whether the detected neutrino comes from a transient point source or has a diffuse origin. This is also corroborated by KM3NeT collaboration which reported that a $2.2\sigma$ upward fluctuation at KM3NeT is required in order to be consistent with the null observation at IceCube KM3NeT:2025npi .
Figure 1: Topography of the vicinity of the KM3NeT and IceCube detectors in the nominal direction of an incoming $\mathcal{O}(100)$ PeV neutrino, as determined by the KM3NeT collaboration. We observe that the path through rock and sea for KM3NeT is an order of magnitude longer than the path through ice en route to the IceCube detector.
In this work, we discuss two scenarios that address this tension. What they have in common is a sterile neutrino. We assume that sterile neutrinos get copiously produced at the source and dominate in flux over active neutrinos. If that were not the case, active neutrino fluxes at the KM3NeT and IceCube detectors would be similar, given the propagation distances in Earth which do not facilitate attenuation of neutrino flux for Standard Model (SM) cross sections. Even if there was a significant attenuation, this would reduce the flux at KM3NeT relative to IceCube, which would further strengthen the tension between the two experiments and is opposite from what we want to achieve. Sterile neutrinos could be produced directly from dark matter annihilation or decay Morrison:2022zwk ; Brdar:2016thq and if they are sufficiently light they will not decay to SM particles before reaching Earth, irrespective of the coupling between SM and the sterile neutrino sector. Another possibility is that the point source is very dense and only weakly interacting particles, such as sterile neutrinos, can exit the production region with significant energy. The lack of multi-messenger observations KM3NeT:2025bxl in association to KM3-230213A event further supports our premise.
The KM3NeT collaboration has very accurately determined the direction of the incoming $\mathcal{O}(100)$ PeV neutrino trajectory, with an uncertainty of $1.5^{\circ}$. The origin of KM3-230213A event corresponds to an incidence angle of $0.6^{\circ}$ at KM3NeT and $\sim 8^{\circ}$ above the horizon at IceCube. Using this, it is straightforward to determine the path length through rock and sea for KM3NeT, as well as the respective path length through ice en route to the IceCube detector for neutrinos coming from the same source. The topography is illustrated in Fig. 1, where we show that the path through Earth for KM3NeT is an order of magnitude longer than that for IceCube. We employ this difference in propagation distance to alleviate the tension between IceCube and KM3NeT. Namely, in what follows, we discuss two mechanisms where sterile neutrino flux partially transitions into active neutrino flux inside the Earth. Due to the longer propagation distance inside the Earth, the active neutrino flux at KM3NeT will be significantly higher compared to IceCube, compensating for the smaller effective area and shorter data-taking time, thereby alleviating the tension.
Neutrino production through matter-induced resonance. When the vacuum mixing between active and sterile neutrinos is very small, matter can resonantly amplify sterile-to-active transitions. The MSW resonance Wolfenstein:1977ue ; Mikheyev:1985zog ; Mikheev:1986wj is realized when the matter potential difference for the involved neutrino species, $V$, equals $[\Delta m^{2}/(2E_{\nu})]\cos 2\theta$, where $\Delta m^{2}$ and $\theta$ are the mass-squared difference and vacuum mixing angle, respectively. Taking the values from the SM, matter potential reads $V\equiv V_{\text{CC}}=\sqrt{2}G_{F}n_{e}\simeq 3.7\times 10^{-23}$ GeV for matter with a density of $1\,\text{g}/\text{cm}^{3}$. The resonant oscillation length then reads $L_{\text{res}}\simeq\pi/(V\sin 2\theta)\approx 5\times 10^{3}\,\theta^{-1}$ km in the limit of small $\theta$. For keV-scale sterile neutrino, which will turn out to be required in our analysis, maximum mixing angle value of $\theta^{2}\simeq 10^{-3}$ is allowed in light of terrestrial experiments MINOS:2017cae ; Bolton:2019pcu (such values are in tension with cosmological constraints, but those can be alleviated Farzan:2019yvo ; Cherry:2016jol ; Chu:2018gxk ). This yields $L_{\text{res}}\simeq 10^{5}$ km or longer, and for $L\approx 150$ km, which corresponds to KM3NeT, the oscillation probability would then not surpass $L^{2}/L_{\text{res}}^{2}\approx 10^{-6}$.
This shortcoming can be addressed by considering a large sterile neutrino matter potential, a scenario that has been widely explored in the literature Pospelov:2011ha ; Pospelov:2012gm ; Kopp:2014fha ; Denton:2018dqq ; Berryman:2018jxt . Specifically, we will focus on the realization introduced in Pospelov:2011ha , where a light vector boson, $V^{\mu}$, associated with a spontaneously broken $U(1)_{B}$ symmetry, interacts with a sterile neutrino, $\nu_{s}$, and quarks Pospelov:2011ha
[…]
Here, $g_{b}^{\prime}$ is a product of the gauge coupling $g_{b}$ and a sterile neutrino charge under a new symmetry group. Since we are interested in the matter potential associated with scattering at momentum transfer $Q^{2}=0$, it is convenient to represent the above terms in the effective field theory framework Pospelov:2011ha ; Kopp:2014fha
$\mathcal{L}_{\text{eff}}\supset\frac{G_{B}}{2}\big{[}\bar{\nu}_{s}\gamma_{\mu}(1-\gamma_{5})\nu_{s}][\bar{p}\gamma^{\mu}p+\bar{n}\gamma^{\mu}n\big{]}\,.$ (2)
This expression is obtained after integrating out $V^{\mu}$ (note that the mass of the introduced vector boson, $m_{V}$, always satisfies $m_{V}^{2}\gg Q^{2}=0$). In Eq. 2, $p$ and $n$ represent protons and neutrons, and $G_{B}=g_{b}g_{b}^{\prime}/m_{V}^{2}$. The sterile neutrino matter potential can then be written as $V_{s}=G_{B}n_{\text{nuc}}$ Pospelov:2011ha ; Kopp:2014fha , where $n_{\text{nuc}}=n_{p}+n_{n}$ is the nucleon number density in matter. For $V_{s}\gg G_{F}n_{e}$, the difference between the sterile and active neutrino matter potential, $V$, is approximately equal to $V_{s}$. We parameterize $V_{s}$ using the dimensionless quantity $\epsilon_{ss}=G_{B}/(\sqrt{2}G_{F})$, which represents the ratio of the new baryonic interaction strength to Fermi’s interaction strength. In the literature Pospelov:2011ha ; Harnik:2012ni ; Pospelov:2012gm , $\epsilon_{ss}$ values up to $\mathcal{O}(10^{2}-10^{3})$ have been explored. Upon investigating the effects of $\epsilon_{ss}$ on neutrino oscillations, we find that values $|\epsilon_{ss}|<10^{3}$ typically lead to deviations smaller than a few percent for atmospheric neutrinos. Note that such an interaction strength precisely addresses the $L^{2}/L_{\text{res}}^{2}\ll 1$ problem for KM3NeT and makes the propagation length in Earth comparable to the resonant oscillation length. We should stress that limits on $\epsilon_{ss}$ become more severe if a sterile neutrino mixes with several active neutrinos; constraints arising from solar neutrino data in a scenario where a sterile neutrino mixes with an electron neutrino are particularly restrictive Kopp:2014fha . However, in this work, we consider sterile neutrino mixing only with a muon neutrino. The generation of muon neutrino flux via oscillations is crucial, as muon neutrinos can produce muon tracks at KM3NeT through SM charged current interactions, explaining the recent observation. For $E_{\nu}=220$ PeV and propagation in water/rock ($\rho\approx 1-2.2\,\text{g}/\text{cm}^{3}$, $n_{\text{nuc}}\approx\rho/m_{p}$), MSW resonance occurs for $\sqrt{\Delta m^{2}}\approx m_{s}\simeq 2\times 10^{-1}\sqrt{\epsilon_{ss}}$ keV, where $m_{s}$ is the sterile neutrino mass. For $\epsilon_{ss}\simeq\mathcal{O}(100)$, a keV-scale sterile neutrino will undergo resonance in the Earth en route to the KM3NeT detector. We do not assume constant density matter, as propagation of sterile neutrinos in the direction corresponding to the detected event at KM3NeT involves passage through both rock and water (see Fig. 1); we take the former (latter) to be $100$ km ($47$ km) long. To this end, we solve the Schrödinger equation which is governed by the following Hamiltonian in the ($\nu_{\mu}$, $\nu_{s}$) flavor basis
$H=\!\begin{pmatrix}c_{\theta}&s_{\theta}\\ -s_{\theta}&c_{\theta}\end{pmatrix}\begin{pmatrix}0&0\\ 0&\frac{m_{s}^{2}}{2E_{\nu}}\end{pmatrix}\begin{pmatrix}c_{\theta}&-s_{\theta}\\ s_{\theta}&c_{\theta}\end{pmatrix}\!+\!\begin{pmatrix}V_{\text{NC}}&0\\ 0&V_{s}\end{pmatrix},$ (3)
where we have abbreviated $c_{\theta}=\cos\theta$ and $s_{\theta}=\sin\theta$. The term $V_{\text{NC}}=-G_{F}n_{n}/\sqrt{2}\approx-V_{\text{CC}}/2$ is the SM neutral current matter potential which, while included in our analysis, will not impact results due to its smallness with respect to $V_{s}=\sqrt{2}G_{F}\epsilon_{ss}(n_{n}+n_{p})$. The resonance occurs in the neutrino sector for negative values of $\epsilon_{ss}$ and this is a scenario that we will adopt (for positive values of $\epsilon_{ss}$ resonance would be observed in the antineutrino sector).
Figure 2: Sterile-to-active neutrino oscillation probability, $P_{s\mu}$, is shown in the parameter space of the sterile neutrino mass $m_{s}$ and interaction strength $-\epsilon_{ss}$ for both KM3NeT and IceCube. The vacuum mixing angle is fixed to $\theta=10^{-2}$. Deeper shades of blue and red correspond to higher oscillation probabilities at KM3NeT and IceCube, respectively.
Figure 3: Sterile-to-active neutrino oscillation probability, $P_{s\mu}$, is shown as a function of interaction strength $-\epsilon_{ss}$ for both KM3NeT and IceCube, for a sterile neutrino mass of $m_{s}=3$ keV and a vacuum mixing angle $\theta=10^{-2}$.
The results of flavor evolution in the Earth for both IceCube and KM3NeT are shown in Figs. 2 and 3; in both figures, the vacuum mixing angle is fixed to $\theta=10^{-2}$ and we present oscillation probabilities after traversing 147 km and 14 km for KM3NeT and IceCube, respectively. In Fig. 2, we show sterile-to-active neutrino oscillation probability, $P_{s\mu}$, in the parameter space of sterile neutrino mass (scanned between $1$ and $10$ keV, values for which a resonance is expected) and $-\epsilon_{ss}$. For KM3NeT, sizable values of $P_{s\mu}$ already appear for $-\epsilon_{ss}\gtrsim 30$. In the region between $-\epsilon_{ss}\simeq 100$ and $-\epsilon_{ss}\simeq 10^{3}$, $P_{s\mu}$ for KM3NeT attains values of $P_{s\mu}\gtrsim 10^{-2}$, implying a relatively large muon neutrino flux generated in Earth. At around $-\epsilon_{ss}\sim 10^{3}$, the oscillation probability for IceCube reaches $P_{s\mu}\simeq 10^{-3}$, remaining smaller by 11112222 orders of magnitude compared to that of KM3NeT. This is because $P_{s\mu}$ depends on $\sin^{2}(L/L_{\text{res}})$, and the neutrino propagation length in Earth (along the direction of the 220 PeV neutrino detected by KM3NeT) is about 10 times shorter at IceCube. Therefore, at resonance, $P_{s\mu}$ is always expected to be at least two orders of magnitude larger for KM3NeT. The figure can be extended to larger values of $-\epsilon_{ss}$ that would require heavier sterile neutrinos to induce a resonance. However, increasing the interaction strength would create tension with the data Pospelov:2011ha ; Harnik:2012ni ; Pospelov:2012gm . In Fig. 3, we show $P_{s\mu}$ as a function of $-\epsilon_{ss}$ for both KM3NeT and IceCube and for a fixed sterile neutrino mass of $m_{s}=3$ keV. The resonance is obtained for $-\epsilon_{ss}\simeq 150$, a parameter point where $P_{s\mu}$ for KM3NeT is two orders of magnitude larger than for IceCube. This difference in oscillation probability, i.e., in the generated muon neutrino flux, alleviates the tension between the two experiments by compensating for KM3NeT’s smaller effective area and shorter data-taking period. In Fig. 3, we also show the dotted line, which represents the probability for sterile-to-active oscillations in vacuum, i.e. the muon neutrino flux generated between the source and the Earth relative to the sterile neutrino flux. The muon neutrino flux generated in vacuum is subdominant compared to the one generated through resonance, as required, since otherwise the event rate would be dominated by neutrinos produced outside the Earth, which would not explain the tension.
Before closing this section, we should also stress that the air density at the KM3NeT location (sea level) is 20% larger than the one at IceCube atmos2 ; atmos1 . Therefore, in principle, a resonance in the atmosphere can be achieved at KM3NeT and avoided at IceCube (see Sponsler:2024iej for resonant neutrino flavor conversion in the atmosphere). However, since rock and water have densities three orders of magnitude larger than air at sea level, the oscillation length at resonance in air would be significantly longer, and we do not anticipate strong oscillation effects in such a case.
Neutrino production via non-standard interactions. In what follows, we present an oscillatory realization that alleviates the tension between IceCube and KM3NeT without involving resonance. We consider the following Hamiltonian in the ($\nu_{\mu},\nu_{s}$) flavor basis
$H=$ $\begin{pmatrix}c_{\theta}&s_{\theta}\\ -s_{\theta}&c_{\theta}\end{pmatrix}\begin{pmatrix}0&0\\ 0&{m_{s}^{2}}/{(2E_{\nu})}\end{pmatrix}\begin{pmatrix}c_{\theta}&-s_{\theta}\\ s_{\theta}&c_{\theta}\end{pmatrix}$
$+\begin{pmatrix}V_{\text{NC}}&\epsilon_{\mu s}V_{\text{CC}}\\ \epsilon_{\mu s}V_{\text{CC}}&0\end{pmatrix}\,,$ (4)
where we now turn on the off-diagonal element in the matter potential. This type of lepton flavor violating term has been constrained in the literature for the active neutrino sector Farzan:2017xzy ; Esteban:2018ppq ; Proceedings:2019qno ; Coloma:2023ixt . Overall, $\mathcal{O}(10^{-1}-1)$ values of $\epsilon_{e\mu}$ remain unconstrained Esteban:2018ppq ; Coloma:2023ixt , and such large non-standard interactions can be generated in UV-complete scenarios Farzan:2015hkd ; Farzan:2016wym ; Farzan:2019xor . As can be seen from Eq. 4, we have an off-diagonal element mixing the muon and sterile neutrino. This term, parameterized by $\epsilon_{\mu s}$, arises from the following effective Lagrangian
$\mathcal{L}\supset-2\sqrt{2}G_{F}\epsilon_{\mu s}^{f}\,(\bar{\nu}_{s}\gamma^{\mu}P_{L}\nu_{\mu})(\bar{f}\gamma_{\mu}f)\,,$ (5)
where $f$ represents SM charged fermions. We will not focus on model-building aspects of this interaction and will instead examine its phenomenological implications, primarily in explaining the tension between KM3NeT and IceCube. Note that the SM fermions that neutrinos interact with in Eq. 5 do not necessarily need to be protons and neutrons, as in our previously discussed case of baryonic interaction; interaction with electrons is viable as well. The $\epsilon_{\mu s}$ coefficient can generally be a complex number, but for simplicity, we take it to be real and positive.
Figure 4: The sterile-to-active neutrino oscillation probability, $P_{s\mu}$, in the scenario with off-diagonal non-standard neutrino interactions is shown for both KM3NeT and IceCube for $m_{s}=100$ eV.
In our scenario, a sterile neutrino with $E_{\nu}=220$ PeV would mix with the active neutrino sector due to matter effects parameterized by $\epsilon_{\mu s}$. The effective mixing angle in matter in the limit of a tiny vacuum mixing angle $\theta$ and large $m_{s}$ takes the form
$\theta_{m}=\frac{1}{2}\tan^{-1}\left(\frac{4\epsilon_{\mu s}E_{\nu}V_{\text{CC}}}{m_{s}^{2}+E_{\nu}V_{\text{CC}}}\right)\,.$ (6)
In the limit of $V_{\text{CC}}\gg m_{s}^{2}/E_{\nu}$, this further reduces to $\theta_{m}=(1/2)\tan^{-1}(4\epsilon_{\mu s})$. The effective mass squared difference in matter for $\theta\approx 0$ reads
$m_{\text{eff}}^{2}=\sqrt{\left(E_{\nu}V_{\text{CC}}+m_{s}^{2}\right)^{2}+16\epsilon_{\mu s}^{2}E_{\nu}^{2}V_{\text{CC}}^{2}}\,\,.$ (7)
Using these results, the conversion probability from sterile to muon neutrino is given by
$P_{s\mu}=\sin^{2}(2\theta_{m})\sin^{2}\left[{m_{\text{eff}}^{2}L}/{(4E_{\nu})}\right]\,\,,$ (8)
where $L$ is neutrino propagation length in Earth. In Fig. 4, we show the conversion probabilities for KM3NeT and IceCube as a function of $\epsilon_{\mu s}$ for $m_{s}=100\text{ eV}$ and $\theta=10^{-4}$. We have scrutinized the impact of the considered interaction on $P_{ee}$, $P_{\mu e}$, and $P_{\mu\mu}$ at low energies and found that $\epsilon_{\mu s}$ values up to $\mathcal{O}(1-10)$ have negligible effect on neutrino oscillations for $m_{s}\gtrsim 100$ eV. Hence, in the figure, we present oscillation probabilities for $\epsilon_{\mu s}$ values up to $10$. The sterile-to-active oscillation probability for KM3NeT is larger than that for IceCube by roughly two orders of magnitude for all presented values of $\epsilon_{\mu s}$. This can be understood from the oscillation formula in Eq. 8, which features an $L^{2}$ dependence when ${m_{\text{eff}}^{2}L}/{(4E_{\nu})}\lesssim\mathcal{O}(1)$, which is satisfied. Thus, the difference in the propagation length between KM3NeT and IceCube leads to the observed difference in the value of $P_{s\mu}$. While the ratio of $P_{s\mu}$ for KM3NeT and IceCube does not depend on $\epsilon_{\mu s}$, it should be noted that solutions with larger $\epsilon_{\mu s}$ are more favorable, as they lead to a larger generated active neutrino flux at KM3NeT, and thus require smaller incoming flux of sterile neutrinos. As in our previous scenario with resonance, the larger muon neutrino flux generated en route to KM3NeT alleviates the tension arising from the non-observation of $\mathcal{O}(100)$ PeV neutrinos in IceCube.
Summary and Outlook. KM3NeT has recently observed the highest-energy neutrino event exceeding $\mathcal{O}(100)$ PeV, and the lack of observation of neutrinos at such energy at IceCube, which has a larger effective area and a longer data-taking time, is unexpected. We present for the first time a new physics solution to this tension that has recently been quantified to be between $2\sigma$ and $3.5\sigma$. We consider a scenario where most of the flux from a point source in the nominal direction of a detected $\mathcal{O}(100)$ PeV neutrino, arrives at Earth as sterile neutrinos. Muon neutrinos are then generated via sterile-to-active oscillations in Earth en route to the KM3NeT and IceCube detectors. Given the direction of the incoming $\mathcal{O}(100)$ PeV neutrino trajectory, the path through Earth for KM3NeT is an order of magnitude longer than for IceCube. This discrepancy in propagation length yields a higher flux of generated muon neutrinos at KM3NeT, which compensates for KM3NeT’s deficiency in effective area and data-taking time, successfully explaining why IceCube has not detected neutrinos at such high energies. We have presented two scenarios that feature strong oscillations between sterile and muon neutrinos: one in which oscillations are resonantly amplified in matter through the sterile neutrino interaction with baryons, and another featuring off-diagonal non-standard interactions. We would like to stress that oscillations may not be the only option for successfully alleviating the tension between KM3NeT and IceCube. In particular, scattering in Earth may be considered, as the likelihood for interaction is proportional to the propagation length in Earth, setting the stage for resolving the tension. However, we have not found a successful solution to the tension when a sterile neutrino produces a muon neutrino via scattering, even in the framework where very large Glashow-like cross sections at $\mathcal{O}(100)$ PeV energies are employed. This is due to the excess of sterile neutrino scattering events within the fiducial volume of IceCube and KM3NeT. We leave further investigation of the scattering method for future work, where we will focus on extended hidden sectors.
Acknowledgments. The work of VB is supported by the United States Department of Energy Grant No. DE-SC0025477.